OpenStudy (anonymous):

Rationalize the denominator of sqrt(-49)/(7-2i)-(4+9i) -49+2i/-40 49+21i/58 77+21i/-2 -77++21i/130 will reward medal

OpenStudy (anonymous):

\[\frac{ \sqrt{-49} }{ (7-2 i)-\left( 4+9\iota \right) }=\frac{ \sqrt{49 \iota ^{2}} }{(7-2\iota)-\left( 4+9\iota \right) }\] \[=\frac{ 7\iota }{ 7-2\iota-4-9\iota }=\frac{ 7\iota }{ 3-11\iota }*\frac{ 3+11\iota }{ 3+11\iota }\] solve it and get the answer.